A Complete Bibliography of Publications in the SIAM Journal on Mathematics of Data Science

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 October 2023 Version 1.08

Title word cross-reference

2r [BNZ21]. 3 [BKSU22]. ℓ_{0} [DHP22]. k [SGN22, Tri19]. q [Rii23]. t [LS19].

-Connectivity [Rii23]. -NN [Tri19]. -SNE [LS19]. -Variance [SGN22].

19 [HNP22].

Abundance [WT22]. Accelerated [BBG20, CC22, DL22, LLP23]. Activation [CP20]. Activations [Ver20]. Active [CZD^{+}20]. Acyclic [LLM22]. Adaptive [HS22, LLRS23]. Adaptivity [HLRJ22]. Admissibility [BNS21]. ADMM [RSC^{+}22]. Adversarial [BGS23, FWDJ23, JC23, LKA^{+}21, WYSO20]. Aggregated [BSV21]. Aggregation [MV22]. Algebraic [BM20]. Algorithm [ANY19, AGL21, CCNP23, ZZST21, ZN22]. Algorithmic [JCD23, LWL20]. Algorithms [AG21, ABR19, BGS23, CDH^{+}21, GPWS21, GCPT20, GH21, JB22, LLRS23, LST19, MWW22, MSK20, MOPS20, VGWS19, XSC22]. Alignment [LMR^{+}20, PWB^{+}19, RBO21]. Alphabets [BM22]. Alternating [LS23]. Analysis [ANY19, BGJ20, BH22, CMS20, CDH^{+}21, CZD^{+}20, DJRW20, GGWX21, HM21, HNP22, HLZ20, KPR^{+}21, MGHV21, MM23, OWXZ21, Rii23]. Anomalies [HNP22]. AO [RSC^{+}22]. AO-ADMM [RSC^{+}22]. Application [ML19]. Applications [CDH^{+}21, DH20, HFL^{+}22, KTM20, Rii23].

Applied [CC22]. Approach [BSV22, HLZ22, RSC^{+}22]. Approximate [Atc21, BK23]. Approximately [UT19]. Approximation [AC20, AD21, AdE22, BGJ20, BGKP19, CLM^{+}22, DMR21, JS21, LLP23, NGBU23, PS21, SGL^{+}20]. Approximations [ABC19]. Archetypal [OWXZ21]. Area [GPWS21]. Argument [TM20]. armed [SRR21]. Arrangements [BV19]. Artificial [BGJ20]. Assess [WWZM20]. Asymmetric [JCD23]. Asymptotic [HLZ20]. Attack [DHP22]. Attacks [CZD^{+}20]. Autodifferentiable [CSAW22]. Automatic [HLZ22]. Averaged [CP20]. Avoid [SM20].

Backtracking [MOPS20]. Balancing [LMM22]. Bandits [SRR21, Zha22]. Barycenter [CC22, HMZ22]. Barycenters [ABA22]. Based [AG21, BFM^{+}21, HFL^{+}22, LMR^{+}20, MRS20, Tri19]. Bayesian [HLZ22, MNO20]. Belief [WWZM20]. Benefit [XSC22]. Benign [WT22]. between [AD21, DB22]. Beyond [WYSO20]. Bi [HHvT22]. Bi-Invariant [HHvT22]. Bias [SRR21]. Big [UT19]. Binary [KT21, WT22]. Biological [CDH^{+}21]. Biwhitening [LZK22]. Black [BGJ20]. Blind [CR22, SST20]. Block [HPG^{+}22, PHLP19, SST20]. Bootstrap [ZG21]. Bounds [AdE22, ERK^{+}19, MD19, WTD^{+}21]. Bregman [CC22, HPG^{+}22, MOPS20]. Bundle [ZG21]. Butterfly [ZRG23].

Capacity [Ver20]. Case [CMR^{+}21, HNP22, KT21]. Categorical [BH22, TLFS21]. Causal [CDS23]. Causality [TLFS21]. Centerpoints [DH20]. Central [dBSLNW23]. Centralities [Ben19]. Centrality [TCTER21]. Certificates [CP20]. Chain [Atc21]. Chains [BFM^{+}21, GCPT20]. Characteristic [GPWS21]. Circular [AGL21, CCNP23]. Classification [DHP22, GPWS21, WT22]. Closing [LPH^{+}23]. Clouds [Tri19]. Cluster [ACQ23, BRZ22]. Clustered [SGN22]. Clustering [BDS21, CS19, CEH23, DM22, LM20, LS19, VGWS19]. Coding [ASE19, LR19]. Coefficients [FT20]. Cohomology [CW20]. Combinations [JB22]. Combined [ZYWL20]. Communication [CBD21, LLC22]. Communication-Efficient [CBD21, LLC22]. Communications [GIMA21]. Communities [ZLZ20]. Community [FKZ22, WWZM20]. Compact [HLZ22]. Comparisons [LST19]. Complete [CG19]. Completion [HZ21, OPAB^{+}21, TTV19]. Complex [CLM^{+}22]. Complex-Valued [CLM^{+}22]. Complexity [ACU23, BH22, GGWX21, PWB^{+}19, RBO21]. Component [CMS20, KT21]. Components [SBCG^{+}19]. Composite [SBCG^{+}19]. Compressive [LM20, ML19]. Computation [HMZ22, LLC22, LMXZ22]. Computation-Optimal [LLC22]. Computational [LST19]. Compute [ABA22]. Concave [MOPS20]. Conditional [KSS20]. Conditioned [BNZ21]. Conditions [JC23]. Connected [BGKP19]. Connectivity [Rii23]. Considerations [ADN^{+}23]. Consistency [Cal19, OWXZ21, SRR21]. Constrained [BTd22, VGWS19]. Constraining [RSC^{+}22]. Constraint [AU21, TLSV20]. Constraints [BLLR22]. Construction [BLLR22]. Contextual [Zha22]. Continuous [MRS20]. Continuously [GPWS21]. Controlled [BK23, CLT20]. Controls [GGWX21]. Convergence [BTd22, BCP22, GGWX21, GH21, LS23, LWL20, TM20, dBSLNW23]. Convergent [ZZST21]. Convex [AAR20, CS19, MOPS20, SNS21, TMTS23, TLFS21]. Convex-Concave [MOPS20].

Convolutional [HH23, LPH^{+}23, ML19]. Convolutions [JB22]. Cooperative [GGWX21]. Coordinate [ADN^{+}23]. Core [TH23, TH19]. Core-Periphery [TH23]. Correct [AG21]. Correlated [LMR^{+}20]. Costs [dBSLNW23]. Count [LZK22]. COVID [HNP22]. COVID-19 [HNP22]. Crystallographic [BE20]. CUR [GH22]. Curse [BGJ20]. Curve [GPWS21]. Cut [LM20].

D [BKSU22]. Data [ACL19a, ACL19b, Bac21, BSV22, Cal19, CW20, CMR^{+}21, DH20, DYM^{+}22, ERK^{+}19, GPWS21, HNP22, HYFB23, Kol19, LKA^{+}21, LMM22, PS21, Sau22, SBCG^{+}19, SGL^{+}20, TMTS23, Tri19, UT19, VFMS23, ZQW^{+}19]. Data-Dependent [ZQW^{+}19]. Data-Driven [TMTS23]. Decentralized [LLC22]. Decoding [ERK^{+}19]. Decomposition [CR22, GH22, KH20, KT21, Sau22]. Decompositions [ARB20, MSK20]. Deconvolution [CR22, KZLW20]. Deep [AD21, AAR20, BFT23, BGJ20, BGK22, BGKP19, CLT20, EGT^{+}21, GRS^{+}20, LPH^{+}23, MD19, NGBU23, PN22]. Defense [WYSO20]. definite [FADS22]. Degree [Cho20]. Denoising [Lee21]. Density [LC23, LMM22]. Dependent [CMR^{+}21, KPR^{+}21, ZQW^{+}19]. Depths [MFM22]. Descent [BHX20, DB22, SM20, TMTS23]. Design [BM22]. Designs [LMR^{+}20]. DESTRESS [LLC22]. Detectability [WWZM20]. Detecting [ZLZ20]. Detection [BRZ22, FKZ22, TH19, TH23]. Determinantal [FSS22]. Determination [HLZ22]. Deterministic [HZ21]. Dictionary [CG19]. Difference [DMR21, KPR^{+}21]. Differential [BGJ20, MFM22]. Differentially [BSV22]. Differentiation [BCP22]. Diffuse [BSV21]. Diffusion [BK23, CDH^{+}21, HTR^{+}23, SBCG^{+}19]. Dimension [BLLR22, BBG20, HS22]. Dimensional [GKBD19, HYFB23, LR19, LMM22, PS21, VFMS23]. Dimensionality [BGJ20, FADS22, HHV22]. Dimensions [Atc21, BBP23, RBO21]. Diploid [BKSU22]. Directed [BFM^{+}21, LLM22, Rii23]. Discovery [DYM^{+}22]. Discrete [BFT23]. Discretizations [BGK22]. Dissimilarity [HHvT22]. Distance [ANY19, BKSU22, DJRW20]. Distances [CDH^{+}21, FILN20, LMM22]. Distributed [CBD21, DMR21, GIMA21]. Distribution [MV22, TLFS21]. Distributional [BH22]. Distributions [CLLD22, HHvT22, HN19]. Divergence [GH21]. Divergences [BDK^{+}21]. Diversity [ABR19, FSS21]. Do [PN22]. Domains [AC20]. Double [BHX20, DB22]. Doubly [LCK21, LC23]. Driven [CP20, TMTS23]. Dual [CC22]. Dynamical [SST20, WTD^{+}21]. Dynamics [CMR^{+}21, LMXZ22].

Effectiveness [Bac21]. Efficient

[ABR19, BGS23, BNZ21, CBD21, LLC22, NRHvBW21, SNS21, ZZST21, ZRG23]. Eigenspace [CBD21]. Eigenvalues [TCTER21]. Eigenvector [Ben19]. Embedding [FADS22]. Embeddings [KSS20]. Empirical [BGJ20, KTM20]. End [HLZ22]. End-to-End [HLZ22]. EnResNet [WYSO20]. Ensemble [CSAW22, WYSO20]. Ensembles [LWL20]. Entries [BNZ21]. Entropic [dBSLNW23]. Entropy [DDD^{+}21]. Equations [BGJ20, MFM22]. Ergodicity [FT20]. Erratum [ACL19a]. Error [BGJ20, Li21, MD19, WTD^{+}21, XCRW22]. Errors [DB22]. Escape [HLZ20]. Estimates [DMS22, LLM22]. Estimation [BM22, BDK^{+}21, CBD21, HN19, WTD^{+}21]. Estimators [BNS21]. Euclidean [BKSU22]. Evaluation [GPWS21, LLP23]. Evolving [GPWS21]. Exact [LMXZ22]. Expanders [HZ21]. Exponential [CLLD22]. Exponential-Wrapped [CLLD22]. Extending [SM20]. Extensions [FADS22]. Extensive [VFMS23]. Extraction [CDJN21, MGHV21]. Extrapolation [Bac21, BTd22, HPG^{+}22]. Extreme [JS21, SNS21].

Factorization [AG21, JCD23]. Factorizations [ZRG23]. FANOK [ARdE21]. Fast [AAR20, BSV21, BRZ22, CDH^{+}21, LS23, ZLRY20, dBSLNW23]. Feature [CDJN21]. Features [BHX20, DYM^{+}22]. Feel [Zha22]. Feel-Good [Zha22]. Few [BNZ21]. Feynman [WYSO20]. Field [GGWX21]. Filters [CSAW22]. Finite [BM22, Cal19, DMR21, HN19, LR19, LLC22]. Finite-Dimensional [LR19]. Finite-Sum [LLC22]. Finite-Time [DMR21]. First [BRZ22]. Fixed [BBP23, Cho20]. Flow [ACQ23]. Forests [LPH^{+}23]. Formalism [WYSO20]. Format [MSK20]. Formulation [DL22]. Framework [MNO20, ZQW^{+}19]. Free [JCD23]. Frequency [LMXZ22]. Fully [CMR^{+}21]. Function [AC20, AD21, DMR21, LLP23]. Functionals [Tri19]. Functions [BV19, JS21, Lee21, NGBU23, PN22]. Fused [ZZST21].

Games [MRS20, YAY23]. GAN [BSCB22]. Gap [AD21, LPH^{+}23]. Gaps [Atc21]. GAT [FWDJ23]. GAT-GMM [FWDJ23]. Gaussian [BLLR22, CR22, DHP22, FWDJ23, GKD19, GKBD19, LCK21, VFMS23, WT22]. General [AC20, EY22]. Generalization [BGJ20, DB22, HHV22, Li21, LMXZ22, XCRW22]. Generalize [OT22]. Generalized [ABN21, GH22]. Generated [HH23]. Generating [BSV22]. Generative [FWDJ23, HH22, JB22, JC23, LKA^{+}21]. Generic [CLT20]. Genome [BKSU22].

Geometric [GKBD19]. Geometry [BKSU22, HTR^{+}23, KZLW20, LC23, LMM22, OT22]. Given [BBG20]. Global [BGS23, Coo21]. GMM [FWDJ23]. GNMR [ZN22]. Good [Zha22]. Gossip [BBG20]. Goursat [SCF^{+}21]. Gradient [AMFF19, ACQ23, CMR^{+}21, HLRJ22, LLRS23, MRS20, MOPS20, SM20]. Gradient-Based [MRS20]. Gradients [KH20]. Granger [TLFS21]. Graph [ABARS20, ANY19, JB22, LMR^{+}20, LLM22, Tri19, TM20, VGWS19]. Graph-Based [LMR^{+}20, Tri19]. Graphical [ZZST21]. Graphs [BSV21, BFM^{+}21, CDS23, Rii23]. Greed [HM21]. Grids [MD19]. Group [SV23]. Group-Invariant [SV23]. Groups [HHvT22]. Guarantee [BDS21]. Guarantees [ABN21, HMZ22, ZZS23].

Hamilton [MFM22]. Hard [ABA22]. Hausdorff [ANY19]. Hawkes [YLB^{+}19]. Hessian [ACU23]. Hessian-Schatten [ACU23]. Heterogeneous [CW20, Lee21]. Heteroscedastic [HYFB23]. Heteroskedastic [LCK21]. Hidden [SBCG^{+}19]. Hierarchical [ABC19]. High [Atc21, HYFB23, LMM22, PS21, RBO21]. High-Dimensional [HYFB23, LMM22, PS21]. Higher [FT20]. Higher-Order [FT20]. Highly [LMR^{+}20]. Hitting [BFM^{+}21]. Holistic [ASE19]. Homology [HNP22, MNO20]. Hypergraph [Ben19, HZ21]. Hypergraphs [CEH23, TH23]. Hyperplane [BV19, DMS22].

Identifiability [AG21, CG19].

Identification

[EH23, GIMA21, SST20, ZRG23].

Identifying [BKSU22]. Ill [BNZ21, TTV19]. Ill-Conditioned [BNZ21]. Ill-Posedness [TTV19]. Image [HHK23]. Images [HH23]. Imaging [HH22]. Immunization [TCTER21]. Impacts [HN19].

Implementation [ADN^{+}23]. Implications [MWW22]. Implicit [FSS21, EGT^{+}21]. Implicitly [FSS22]. Importance [ZLRY20]. Improved [HM21, dBSLNW23]. Improvement [LM20]. Inconsistency [BBP23]. Independent [YAY23]. Induced [GCPT20]. Inequality [BLLR22]. Inertial [MOPS20]. Inference [LC23]. Infinite [Cal19]. Information [CW20, ERK^{+}19]. Information-Theoretic [ERK^{+}19]. Informed [BLG^{+}22]. Inhomogeneous [HTR^{+}23]. Initialized [LS23]. Input [TMTS23]. Input-Convex [TMTS23]. Insights [PN22]. Instance [KPR^{+}21, LLP23]. Instance-Dependent [KPR^{+}21]. Instance-Optimal [LLP23]. Interface [BSV21]. Interpolation [CLT20, Li21, RDB20, XCRW22, XSC22]. Interpretable [PS21]. Intrinsic [FILN20, HS22, ZQW^{+}19]. Introduction [Kol19]. Invariant [ADN^{+}23, HHvT22, SV23]. Inverse [HH22, HFL^{+}22]. ISLET [ZLRY20]. Isometric [HHK23]. Iterations [BBG20]. Iterative [BNZ21].

Jacobi [BBG20, MFM22]. Joining [JAA^{+}21]. Joint [FKZ22]. Jointly [DYM^{+}22]. Journal [Kol19].

Kac [WYSO20]. Kaczmarz [HM21]. Kalman [CSAW22]. Kernel [BBP23, FSS21, HS22, LCK21, Li21, SCF^{+}21]. Kernels [MV22]. Kinds [PN22]. Knockoffs [ARdE21].

Label [JB22]. Labeled [Cal19]. Landmarking [GKD19, GKBD19]. Langevin [CMR^{+}21, LLRS23]. Laplacian [TM20]. Laplacians [MWW22, MM23]. Large [KH20]. Large-Scale [KH20]. Lasso [ABN21, ZZST21]. Latent [JAA^{+}21]. Layer [ASE19, BGS23, GRS^{+}20]. Layer-Parallel [GRS^{+}20]. Layered [CP20]. Learn [PN22].

Learning [ACU23, BK23, BSV21, BH22, Cal19, CDS23, CMS20, CG19, DB22, DMR21, EGT^{+}21, GIMA21, GGWX21, HHK23, JB22, JS21, KPR^{+}21, LR19, LLRS23, MRS20, SV23, YAY23, Zha22, ZQW^{+}19]. Least [ABC19, BNZ21, LS23]. Lie [HHvT22]. Like [NRHvBW21]. Limit [GCPT20, dBSLNW23]. Limits [Tri19]. Line [ZN22]. Linear [ANY19, ARdE21, BM22, DB22, DMR21, LLP23, LMXZ22, ML19, Ver20]. Linear-Time [ANY19]. Linearly [ZZST21]. Lipschitz [Cal19, CP20, NGBU23]. Local [CDS23, DM22, LM20, MV22]. Location [ZYWL20]. Logistic [DH20]. Loss [Lee21]. Low [ARB20, BNZ21, MSK20, SGL^{+}20, UT19, ZLRY20, ZN22]. Low-Rank [ARB20, MSK20, SGL^{+}20, ZLRY20].

Machines [JS21]. Majorization [HPG^{+}22]. Manifold [BM20, FILN20, HHK23, HLZ20, LC23, LKA^{+}21, LPH^{+}23]. Manifolds [DL22, GKD19]. Mapping [HHK23]. Maps [LR19]. Markov [Atc21, BFM^{+}21, GCPT20, MV22]. MARL [GGWX21]. Matching [CMS20]. Mathematical [BE20]. Mathematics [Kol19]. Matrices [BNZ21, UT19]. Matrix [AG21, BSV21, EY22, GH22, JCD23, LZK22, LS23, Lee21, OPAB^{+}21, Sau22, TTV19, ZRG23, ZN22]. Matrix-Vector [BSV21]. Max [JC23]. Maximization [PHLP19]. Maximum [DDD^{+}21, TM20]. MCMC [AMFF19]. Mean [GGWX21, KSS20]. Mean-Field [GGWX21]. Means [SRR21]. Measure [ABR19, MGHV21, MM23]. Measures [HHvT22]. Measuring [ACU23, LWL20]. Memorizing [ZQW^{+}19]. Memory [SNS21, Ver20]. Memory-Efficient [SNS21]. Method [BCP22, BSCB22, TH19]. Methods [BM22, BSV21, CC22, DDD^{+}21, FSS21, HS22, HLRJ22, OPAB^{+}21, ZLZ20].

Metric [BFM^{+}21, MM23, VGWS19]. Metric-Constrained [VGWS19]. Min [ABN21, JC23]. Min-Max [JC23]. Minima [Coo21]. Minimax [BM22]. Minimization [BGJ20, HPG^{+}22]. Minimum [DJRW20, Li21]. Mining [Sau22]. Mirror [TMTS23]. Missing [VFMS23]. Mixing [Atc21]. Mixture [CR22, DHP22, FWDJ23, TLFS21]. Mixtures [HN19, WT22]. Model [CDJN21, DHP22, HH22, JB22, JCD23]. Model-Free [JCD23]. Modelling [VFMS23]. Models [AMFF19, BM22, BHX20, CR22, FWDJ23, JAA^{+}21, LST19, PHLP19, SST20]. Modes [RSC^{+}22]. Modularity [PHLP19, WWZM20]. Moments [Cho20]. Mondrian [OT22]. Morphometrics [GKBD19]. Motzkin [HM21]. Moving [ACQ23]. MR [LKA^{+}21]. MR-AN [LKA^{+}21]. Multi [ASE19, RBO21, SRR21]. Multi-armed [SRR21]. Multi-Layer [ASE19]. Multi-Reference [RBO21]. Multiagent [YAY23]. Multiclass [DH20]. Multicomponent [CW20]. Multigraphs [Cho20]. Multikernel [AU21]. Multilayer [BSV21, PHLP19, WWZM20]. Multilinear [ML19]. Multimodal [DYM^{+}22, SBCG^{+}19]. Multiple [ZZST21]. Multireference [PWB^{+}19]. Multitemporal [CDH^{+}21]. Multivariate [AC20, YLB^{+}19].

Near [AC20, BK23]. Near-Optimal [AC20]. Nearest [XSC22]. Neighbor [JAA^{+}21, XSC22]. Neighborhood [MGHV21]. Network [BDK^{+}21, CP20, RDB20, Rii23, YLB^{+}19]. Networks [AD21, AAR20, BGS23, BFT23, BLG^{+}22, BGJ20, BBG20, BGK22, BGKP19, BRZ22, CLM^{+}22, Coo21, CDH^{+}21, CLT20, CZD^{+}20, GRS^{+}20, HH23, HLZ22, JC23, LKA^{+}21, LLM22, LPH^{+}23, LLRS23, ML19, MD19, NGBU23, NRHvBW21, PHLP19, PN22, SV23, TMTS23, TH19, Ver20, ZLZ20, ZZS23]. Neural [AD21, AAR20, BFT23, BLG^{+}22, BGJ20, BDK^{+}21, BGK22, BGKP19, CLM^{+}22, Coo21, CLT20, CZD^{+}20, GRS^{+}20, HH23, HLZ22, LLM22, LLRS23, NGBU23, NRHvBW21, PN22, TMTS23, Ver20, ZZS23]. NN [Tri19]. Node [TCTER21]. Noise [LCK21, MV22]. Noiseless [BSV22]. Noisy [EY22, FILN20]. Nonbacktracking [CEH23, TCTER21]. Nonconcave [JC23]. Nonconvex [CMR^{+}21, HLRJ22, JC23, LLC22, LLRS23, MOPS20]. Nonconvex-Nonconcave [JC23]. Nonlinear [LR19, LLM22, OPAB^{+}21, Sau22, TH19]. Nonsmooth [JC23]. Nonuniform [CEH23]. Norm [Li21]. Normal [ZG21]. Normal-Bundle [ZG21]. Normalization [LCK21]. Norms [DM22]. Note [HH23]. Notion [SGN22]. NP [ABA22]. NP-Hard [ABA22]. Numerical [ADN^{+}23, BGJ20].

Objects [MM23]. Oblique [LPH^{+}23]. Observations [SST20]. ODEs [CLT20]. One [ZN22]. One-Line [ZN22]. Operating [GPWS21]. Operator [EY22]. Operators [CP20, SBCG^{+}19]. Optimal [AC20, AGL21, ABC19, BGKP19, CCNP23, CC22, KPR^{+}21, KTM20, LLC22, LLP23, ZLRY20, ZYWL20]. Optimality [BK23, JC23]. Optimally [HYFB23]. Optimization [ABN21, BGS23, BNS21, BRZ22, DL22, HLRJ22, HLZ20, LLC22, LST19, MOPS20, SNS21, VGWS19]. Order [BRZ22, CR22, FT20]. Order-3 [CR22]. Organisms [BKSU22]. Organization [BKSU22]. Out-of-Sample [FADS22]. Overcomes [BGJ20]. Overcomplete [CR22]. Overfitting [WT22]. Overlapping [ZLZ20]. Overparameterization [XCRW22]. Overparameterized [Coo21, HHV22]. Overparametrized [JCD23].

Pairs [GH22]. PARAFAC2 [RSC^{+}22].

Parallel [GRS^{+}20]. Parameter [HN19]. Part [KT21]. Partial [BGJ20]. Partitioning [HS22]. Path [LMM22]. Paths [BFT23, YAY23]. Pattern [MGHV21, ZQW^{+}19]. PCA [HYFB23, TTV19]. PDE [SCF^{+}21]. Penalty [ABN21]. Performance [BDS21, DMR21, LST19]. Periphery [TH23, TH19]. Persistent [CW20, HNP22, MNO20, MWW22]. Phase [BE20, RBO21, TLSV20]. PIDE [BGK22]. Piggyback [BCP22]. Piggyback-Style [BCP22]. Point [BCP22, FSS22, SNS21, ZZST21]. Points [BM20, SM20]. Policy [LLP23]. Polynomial [BV19, BBG20]. Pooled [ERK^{+}19]. Posedness [TTV19]. Positive [FADS22, KT21]. Positive-Semidefinite [KT21]. Post [ZZS23]. Post-training [ZZS23]. Powering [ABARS20]. Practical [ABN21]. Practice [AD21, DDD^{+}21]. Presence [VFMS23]. Primal [CC22]. Primal-Dual [CC22]. Principle [LMXZ22, TM20]. Priori [LLM22]. Private [BSV22]. Pro [NRHvBW21]. Probabilistic [VFMS23]. Probabilities [BFM^{+}21]. Problem [BE20, EH23, MGHV21]. Problems [BCP22, CC22, FSS22, HH22, HFL^{+}22, JC23, LMR^{+}20, ZZST21]. Procedure [DJRW20]. Process [GKD19, GKBD19, OT22, VFMS23]. Processes [BLLR22, BK23, FSS22, FT20, YLB^{+}19]. Products [BSV21]. Programming [FKZ22, YTF^{+}21]. Programs [AdE22]. Projection [NRHvBW21]. Projections [HFL^{+}22]. Propagation [JB22, WWZM20]. Properties [GCPT20, JS21, MWW22]. Provable [BLG^{+}22, ZZS23, ZN22]. Provably [AG21, LS19]. Proximal [LST19, MOPS20, ZZST21]. Pruning [AAR20, BLG^{+}22].

Q [BK23, GGWX21]. Q-Learning [GGWX21]. Quadratic [BGS23]. Quadratic-Time [BGS23]. Quantitative [CLM^{+}22]. Quantization [ZZS23].

Random [BV19, BM20, Cho20, DMS22, LPH^{+}23, MM23]. Randomized [HMZ22, LWL20, MSK20]. Randomly [LS23]. Rank [ARB20, BNZ21, HLZ22, LZK22, MSK20, SGL^{+}20, UT19, ZLRY20, ZN22]. Rankability [ACL19a, ACL19b]. Rates [dBSLNW23]. Receiver [GPWS21]. Recommendations [ABR19]. Reconstruction [FILN20, JAA^{+}21, LR19, YLB^{+}19]. Recovering [CS19, SBCG^{+}19]. Recovery [ABN21, BNZ21, ZN22]. Rectified [Ver20]. Recursive [GCPT20, GH21]. Reduction [BLLR22, FADS22, HHV22]. Reference [RBO21]. Registration [VFMS23]. Regression [ARB20, AU21, BBP23, DB22, DH20, FSS22, LMR^{+}20, LWL20, ZLRY20]. Regressions [HHV22]. Regressors [TM20]. Regularity [HH23]. Regularization [FSS21, HHV22, JCD23, LMR^{+}20, WT22, ZQW^{+}19]. Regularize [FSS22]. Regularized [KTM20, LKA^{+}21, ZZST21]. Reinforcement [BH22, YAY23, Zha22]. Relating [PHLP19]. Relative [ANY19]. ReLU [BGS23, MD19]. Removal [TCTER21]. Rényi [BDK^{+}21]. Representations [BDK^{+}21]. Resampling [HMZ22]. Residual [GRS^{+}20]. ResNets [WYSO20]. Restriction [SM20]. Results [CLM^{+}22]. Retrieval [BE20, TLSV20]. Reuse [GPWS21]. Reveals [LZK22]. Richardson [Bac21]. Ridge [JS21]. Ridgeless [BBP23]. Riemannian [DL22, FILN20]. Rigidity [TTV19]. Rigorous [KSS20]. Risk [BGJ20, SRR21]. Robust [DHP22, LCK21, LC23, TTV19]. Robustness [ABARS20]. Role [MGHV21]. Rotational [FKZ22]. Rough [BFT23]. Rules [EH23].

Saddle [BCP22, HLZ20, SM20]. Saddle-Point [BCP22]. Safe [EH23]. Sample [FADS22, HHvT22, PWB^{+}19, RBO21, SRR21]. Sampling [AC20, ABC19, BSV22, FSS21, HM21, MV22, SNS21, Zha22]. Satisficing [YAY23]. Scalable [YTF^{+}21]. Scale [KH20]. Scaling [LC23]. Schatten [ACU23]. Scheme [BTd22]. Schemes [ACU23, LR19]. Scholes [BGJ20]. Science [Bac21, Kol19]. Scientific [LKA^{+}21]. SCOTT [ZYWL20]. Selection [ADN^{+}23, DJRW20]. Semi [BSV21, FADS22]. Semi-definite [FADS22]. Semi-supervised [BSV21]. Semidefinite [FKZ22, KT21, YTF^{+}21]. Semiparametric [FSS22]. Sensing [LM20, LS23, ML19]. Sensitivity [BLG^{+}22]. Sensitivity-Informed [BLG^{+}22]. Separability [DH20]. Sequences [Cho20]. Sequential [ABC19, BLLR22]. Seriation [AGL21, CCNP23]. Series [TLFS21]. Sets [HLZ20]. Setting [LWL20]. Shape [VFMS23, ZYWL20]. Shape-Location [ZYWL20]. Sharp [DMS22, ERK^{+}19]. Shifting [EY22]. Short [KZLW20]. Short-and-Sparse [KZLW20]. SIAM [Kol19]. Signals [Lee21]. Signature [SCF^{+}21]. Similarity [MGHV21]. SIMODS [Kol19]. Simple [CCNP23]. Simplex [AG21]. Simplex-Structured [AG21]. Simplicial [Rii23]. Singularity [HN19]. Size [SM20]. Sketching [ZLRY20]. SLOPE [EH23]. Smooth [DYM^{+}22]. SNE [LS19]. Soft [ABN21]. Soft-Min [ABN21]. Solution [BNS21, SCF^{+}21]. Solutions [BCP22, EH23]. Some [GCPT20]. Space [AMFF19]. Spaces [ABC19, CLLD22]. Sparse [ASE19, ARB20, ABN21, AdE22, GIMA21, KZLW20, MD19, Sau22, TLSV20, ZRG23]. Sparsely [BGKP19]. Sparsity [AG21, AU21]. Sparsity-Based [AG21]. Spatial [HNP22]. Spatiotemporal [HNP22, YLB^{+}19]. Spectral [ABARS20, Atc21, BDS21, CEH23, DYM^{+}22, JAA^{+}21, TH19, WTD^{+}21, ZLZ20]. Speedy [BH22]. Spheres [LST19]. Spline [NGBU23, PN22]. Squares [ABC19, BNZ21, LS23]. Stability [BFT23]. Standard [BCP22]. State [AMFF19, CDH^{+}21]. Statistical [BM22, HMZ22, KTM20, MM23]. Step [SM20]. Step-Size [SM20]. Steps [BSCB22]. Stochastic [AMFF19, BNS21, CMR^{+}21, DH20, FT20, GCPT20, GH21, HLRJ22, KH20, LCK21, LC23, LLRS23, OT22, PHLP19, SST20, YAY23]. Stop [ZQW^{+}19]. Strategies [AC20]. Streaming [SGL^{+}20]. Streams [CMR^{+}21]. Strict [AGL21, CCNP23, HLZ20, SM20]. Stride [BSCB22]. Structural [CDS23, CZD^{+}20]. Structure [WWZM20]. Structured [AG21, SNS21]. Structures [CP20, HN19]. Studies [HNP22]. Style [BCP22]. Subspace [CZD^{+}20]. Sum [DM22, LLC22]. Sum-of-Norms [DM22]. Supervised [SV23, BSV21]. Support [WT22]. Symmetric [CLLD22]. Symmetry [KZLW20]. Synchronization [FKZ22]. Synthesis [DDD^{+}21]. Synthetic [BSV22]. System [ABR19]. System-Wide [ABR19]. Systems [EY22].

Tail [DJRW20]. Taking [BSCB22]. Taming [LLRS23]. Targeted [TCTER21]. Tasks [DB22]. Temporal [DMR21, KPR^{+}21]. Temporal-Difference [DMR21]. Tensor [ARB20, CR22, HZ21, HLZ22, KH20, MSK20, OPAB^{+}21, SV23, SGL^{+}20, ZLRY20]. Tensors [BV19]. Tessellations [DMS22]. Test [GPWS21]. Texture [DDD^{+}21]. Theorem [dBSLNW23]. Theorems [DH20]. Theoretic [ERK^{+}19]. Theory [AD21, BE20, DDD^{+}21, KTM20, KSS20, PN22]. Thompson [Zha22]. Three [Ben19, GKBD19, VFMS23]. Three-Dimensional [GKBD19, VFMS23]. Threshold [BV19, DJRW20, Ver20]. Time [ANY19, ARdE21, BGS23, BSCB22, DMR21, HTR^{+}23, TLFS21]. Time-Inhomogeneous [HTR^{+}23]. Times [Atc21]. Topology [HTR^{+}23]. Total [ACU23]. Tracking [ZYWL20]. Train [NRHvBW21, SV23]. Training [BGS23, BSCB22, FWDJ23, GRS^{+}20, HLZ22, NRHvBW21, ZZS23]. Transfer [CMS20, DB22]. Transition [RBO21]. Transport [CC22, CDJN21, KTM20, ZYWL20]. Transportation [dBSLNW23]. Tree [ACQ23, JAA^{+}21]. Trees [CS19]. Trigonometric [XCRW22]. Trimmed [ABN21]. Tseng [BSCB22]. Tucker [ARB20, MSK20, SGL^{+}20]. Tukey [MFM22]. TUSLA [LLRS23]. Tverberg [DH20]. Two [BGS23, BHX20, BSCB22]. Two-Layer [BGS23].

Uniform [Cho20, TM20]. Unifying [JB22]. Unit [Ver20]. Universal [CLT20, CZD^{+}20]. Unknown [BM22]. Unlabeled [Cal19]. Using [ARB20, ACU23, BSV21, BBG20, GPWS21, GH21, HNP22, MD19, NGBU23, ZLZ20].

Valued [CLM^{+}22]. Vanishing [MV22]. Var [NRHvBW21]. Variable [NRHvBW21]. Variance [SGN22]. Variation [ACU23]. Variational [BDK^{+}21, DL22, HH22, PN22, Tri19]. Vector [BTd22, BSV21]. Vectors [WT22]. Versions [DM22]. Via [MM23, BFT23, BKSU22, CDS23, FKZ22, HLZ22, LLRS23, SNS21, WYSO20, ZLRY20].

Wasserstein [ABA22, CC22, GH21, HHK23, HFL^{+}22, HMZ22]. Wasserstein-Based [HFL^{+}22]. Wassmap [HHK23]. Way [ASE19]. Weak [BHX20]. Weighted [ABC19, HYFB23, Lee21, Li21, LLM22, XCRW22]. Wide [ABR19]. Works [HM21]. Wrapped [CLLD22].

X [TCTER21]. X-Centrality [TCTER21].

Zeros [EH23, Sau22].

References

Aghasi:2020:FCP

[AAR20] Alireza Aghasi, Afshin Abdi, and Justin Romberg. Fast convex pruning of deep neural networks. SIAM Journal on Mathematics of Data Science, 2(1): 158–188, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Altschuler:2022:WBN

[ABA22] Jason M. Altschuler and Enric Boix-Adserà. Wasserstein barycenters are NP-hard to compute. SIAM Journal on Mathematics of Data Science, 4(1):179–203, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                                    //epubs.siam.org/doi/10.
                                                    1137/21M1390062.

Abbe:2020:GPS

[ABARS20] Emmanuel Abbe, Enric Boix-Adserà, Peter Ralli, and Colin Sandon. Graph powering and spectral robustness. SIAM Journal on Mathematics of Data Science, 2(1):132–157, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Arras:2019:SSO

[ABC19] Benjamin Arras, Markus Bachmayr, and Albert Cohen. Sequential sampling for optimal weighted least squares approximations in hierarchical spaces.

SIAM Journal on Mathematics of Data Science, 1(1):189–207, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Amir:2021:TLS

[ABN21] Tal Amir, Ronen Basri, and Boaz Nadler. The trimmed lasso: Sparse recovery guarantees and practical optimization by the generalized softmin penalty. SIAM Journal on Mathematics of Data Science, 3(3):900–929, ???? 2021. CO-DEN SJMDAQ. ISSN 2577-0187.

Antikacioglu:2019:NSW

[ABR19] Arda Antikacioglu, Tanvi Bajpai, and R. Ravi. A new system-wide diversity measure for recommendations with efficient algorithms. SIAM Journal on Mathematics of Data Science, 1(4):759–779, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Adcock:2020:NOS

[AC20] Ben Adcock and Juan M. Cardenas. Near-optimal sampling strategies for multivariate function approximation on general domains. SIAM Journal on Mathematics of Data Science, 2(3):607–630, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Anderson:2019:ERD

[ACL19a] Paul Anderson, Timothy Chartier, and Amy Langville. Erratum: The rankability of data. SIAM Journal on Mathematics of Data Science, 1(3): 643–646, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187. See [ACL19b].

Anderson:2019:RD

[ACL19b] Paul Anderson, Timothy Chartier, and Amy Langville. The rankability of data. SIAM Journal on Mathematics of Data Science, 1(1):121–143, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187. See erratum [ACL19a].

Arias-Castro:2023:MCT

[ACQ23] Ery Arias-Castro and Wanli Qiao. Moving up the cluster tree with the gradient flow. SIAM Journal on Mathematics of Data Science, 5(2):400–421, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1469869.

Aziznejad:2023:MCL

[ACU23] Shayan Aziznejad, Joaquim Campos, and Michael Unser. Measuring complexity of learning schemes using Hessianschatten total variation. SIAM Journal on Mathematics of Data Science, 5(2):422–445, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M147517X.

Adcock:2021:GBT

[AD21] Ben Adcock and Nick Dexter. The gap between theory and practice in function approximation with deep neural networks. SIAM Journal on Mathematics of Data Science, 3(2): 624–655, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Askari:2022:ABS

[AdE22] Armin Askari, Alexandre d’Aspremont, and Laurent El Ghaoui. Approximation bounds for sparse programs. SIAM Journal on Mathematics of Data Science, 4(2):514–530, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1398677.

Archimbaud:2023:NCN

[ADN^{+}23] Aurore Archimbaud, Zlatko Drmac, Klaus Nordhausen, Una Radojicić, and Anne Ruiz-Gazen. Numerical considerations and a new implementation for invariant coordinate selection. SIAM Journal on Mathematics of Data Science, 5(1):97–121, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1498759.

Abdolali:2021:SSM

[AG21] Maryam Abdolali and Nicolas Gillis. Simplex-structured matrix factorization: Sparsitybased identifiability and provably correct algorithms. SIAM Journal on Mathematics of Data Science, 3(2):593–623, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Armstrong:2021:OAS

[AGL21] Santiago Armstrong, Cristóbal Guzmán, and Carlos A. Sing Long. An optimal algorithm for strict circular seriation. SIAM Journal on Mathematics of Data Science, 3(4):1223– 1250, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Aicher:2019:SGM

[AMFF19] Christopher Aicher, Yi-An Ma, Nicholas J. Foti, and Emily B. Fox. Stochastic gradient MCMC for state space models. SIAM Journal on Mathematics of Data Science, 1(3):555–587, ???? 2019. CO-DEN SJMDAQ. ISSN 2577-0187.

Aksoy:2019:LTA

[ANY19] Sinan G. Aksoy, Kathleen Nowak, and Stephen J. Young. A linear-time algorithm and analysis of graph relative Hausdorff distance. SIAM Journal on Mathematics of Data Science, 1(4):647–666, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Ahmed:2020:TRU

[ARB20] Talal Ahmed, Haroon Raja, and Waheed U. Bajwa. Tensor regression using low-rank and sparse Tucker decompositions. SIAM Journal on Mathematics of Data Science, 2(4):944–966, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Askari:2021:FKL

[ARdE21] Armin Askari, Quentin Rebjock, Alexandre d’Aspremont, and Laurent El Ghaoui. FANOK: Knockoffs in linear time. SIAM Journal on Mathematics of Data Science, 3(3): 833–853, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Aberdam:2019:MLS

[ASE19] Aviad Aberdam, Jeremias Sulam, and Michael Elad. Multilayer sparse coding: The holistic way. SIAM Journal on Mathematics of Data Science, 1(1):46–77, ???? 2019. CO-DEN SJMDAQ. ISSN 2577-0187.

Atchade:2021:ASG

[Atc21] Yves F. Atchadé. Approximate spectral gaps for Markov chain mixing times in high dimensions. SIAM Journal on Mathematics of Data Science, 3(3): 854–872, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Aziznejad:2021:MRS

[AU21] Shayan Aziznejad and Michael Unser. Multikernel regression with sparsity constraint. SIAM Journal on Mathematics of Data Science, 3(1):201–224, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Bach:2021:ERE

[Bac21] Francis Bach. On the effectiveness of Richardson extrapolation in data science.

SIAM Journal on Mathematics of Data Science, 3(4):1251– 1277, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Berthier:2020:AGN

[BBG20] Raphaël Berthier, Francis Bach, and Pierre Gaillard. Accelerated gossip in networks of given dimension using Jacobi polynomial iterations. SIAM Journal on Mathematics of Data Science, 2(1):24–47, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Beaglehole:2023:IKR

[BBP23] Daniel Beaglehole, Mikhail Belkin, and Parthe Pandit. On the inconsistency of kernel ridgeless regression in fixed dimensions. SIAM Journal on Mathematics of Data Science, 5(4):854–872, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1499819.

Bogensperger:2022:CPS

[BCP22] Lea Bogensperger, Antonin Chambolle, and Thomas Pock. Convergence of a piggybackstyle method for the differentiation of solutions of standard saddle-point problems. SIAM Journal on Mathematics of Data Science, 4(3):1003–1030, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1455887.

Birrell:2021:VRN

[BDK^{+}21] Jeremiah Birrell, Paul Dupuis, Markos A. Katsoulakis, Luc Rey-Bellet, and Jie Wang. Variational representations and neural network estimation of Rényi divergences. SIAM Journal on Mathematics of Data Science, 3(4):1093–1116, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Boedihardjo:2021:PGS

[BDS21] March Boedihardjo, Shaofeng Deng, and Thomas Strohmer. A performance guarantee for spectral clustering. SIAM Journal on Mathematics of Data Science, 3(1):369–387, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Bendory:2020:TMT

[BE20] Tamir Bendory and Dan Edidin. Toward a mathematical theory of the crystallographic phase retrieval problem. SIAM Journal on Mathematics of Data Science, 2(3): 809–839, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Benson:2019:THE

[Ben19] Austin R. Benson. Three hypergraph eigenvector centralities. SIAM Journal on Mathematics of Data Science, 1(2): 293–312, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Boyd:2021:MDG

[BFM^{+}21] Zachary M. Boyd, Nicolas Fraiman, Jeremy Marzuola, Peter J. Mucha, Braxton Osting, and Jonathan Weare. A metric on directed graphs and Markov chains based on hitting probabilities. SIAM Journal on Mathematics of Data Science, 3(2):467–493, ???? 2021. CO-DEN SJMDAQ. ISSN 2577-0187.

Bayer:2023:SDN

[BFT23] Christian Bayer, Peter K. Friz, and Nikolas Tapia. Stability of deep neural networks via discrete rough paths. SIAM Journal on Mathematics of Data Science, 5(1):50–76, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1472358.

Berner:2020:AGE

[BGJ20] Julius Berner, Philipp Grohs, and Arnulf Jentzen. Analysis of the generalization error: Empirical risk minimization over deep artificial neural networks overcomes the curse of dimensionality in the numerical approximation of Black–Scholes partial differential equations. SIAM Journal on Mathematics of Data Science, 2(3):631–657, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Bohn:2022:DNN

[BGK22] Bastian Bohn, Michael Griebel, and Dinesh Kannan. Deep neural networks and PIDE discretizations. SIAM Journal on Mathematics of Data Science, 4(3):1145–1170, ????

2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1438554.

Bolcskei:2019:OAS

[BGKP19] Helmut Bölcskei, Philipp Grohs, Gitta Kutyniok, and Philipp Petersen. Optimal approximation with sparsely connected deep neural networks. SIAM Journal on Mathematics of Data Science, 1(1):8–45, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Bai:2023:EGO

[BGS23] Yatong Bai, Tanmay Gautam, and Somayeh Sojoudi. Efficient global optimization of two-layer ReLU networks: Quadratic-time algorithms and adversarial training. SIAM Journal on Mathematics of Data Science, 5(2):446–474, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1467134.

Bock:2022:SCD

[BH22] Markus Böck and Clemens Heitzinger. Speedy categorical distributional reinforcement learning and complexity analysis. SIAM Journal on Mathematics of Data Science, 4(2):675–693, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/20M1364436.

Belkin:2020:TMD

[BHX20] Mikhail Belkin, Daniel Hsu, and Ji Xu. Two models of double descent for weak features. SIAM Journal on Mathematics of Data Science, 2(4):1167– 1180, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Bayraktar:2023:AQL

[BK23] Erhan Bayraktar and Ali Devran Kara. Approximate Q learning for controlled diffusion processes and its near optimality. SIAM Journal on Mathematics of Data Science, 5(3):615–638, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1484201.

Belyaeva:2022:IGO

[BKSU22] Anastasiya Belyaeva, Kaie Kubjas, Lawrence J. Sun, and Caroline Uhler. Identifying 3D genome organization in diploid organisms via Euclidean distance geometry. SIAM Journal on Mathematics of Data Science, 4(1):204–228, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1390372.

Baykal:2022:SIP

[BLG^{+}22] Cenk Baykal, Lucas Liebenwein, Igor Gilitschenski, Dan Feldman, and Daniela Rus. Sensitivity-informed provable pruning of neural networks. SIAM Journal on Mathematics of Data Science, 4(1):26–45, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/20M1383239.

Bachoc:2022:SCD

[BLLR22] François Bachoc, Andrés F.

López-Lopera, and Olivier Roustant. Sequential construction and dimension reduction of Gaussian processes under inequality constraints. SIAM Journal on Mathematics of Data Science, 4(2):772–800, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1407513.

Breiding:2020:RPA

[BM20] Paul Breiding and Orlando Marigliano. Random points on an algebraic manifold. SIAM Journal on Mathematics of Data Science, 2(3):683–704, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Behr:2022:SMM

[BM22] Merle Behr and Axel Munk.

Statistical methods for minimax estimation in linear models with unknown design over finite alphabets. SIAM Journal on Mathematics of Data Science, 4(2):490–513, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1398860.

Basu:2021:ASE

[BNS21] Amitabh Basu, Tu Nguyen, and Ao Sun. Admissibility of solution estimators for stochastic optimization. SIAM Journal on Mathematics of Data Science, 3(1):31–51, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Bauch:2021:RIL

[BNZ21] Jonathan Bauch, Boaz Nadler, and Pini Zilber. Rank 2r iterative least squares: Efficient recovery of ill-conditioned low rank matrices from few entries. SIAM Journal on Mathematics of Data Science, 3(1):439–465, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Bomze:2022:FCD

[BRZ22] Immanuel M. Bomze, Francesco Rinaldi, and Damiano Zeffiro. Fast cluster detection in networks by first order optimization. SIAM Journal on Mathematics of Data Science, 4(1):285–305, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1408658.

Bohm:2022:TST

[BSCB22] Axel Böhm, Michael Sedlmayer, Ernö Robert Csetnek, and Radu Ioan Bot. Two steps at a time — taking GAN training in stride with Tseng’s method. SIAM Journal on Mathematics of Data Science, 4(2):750–771, ????

2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1420939.

Bergermann:2021:SSL

[BSV21] Kai Bergermann, Martin Stoll, and Toni Volkmer. Semisupervised learning for aggregated multilayer graphs using diffuse interface methods and fast matrix-vector products. SIAM Journal on Mathematics of Data Science, 3(2): 758–785, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Boedihardjo:2022:PSN

[BSV22] March Boedihardjo, Thomas Strohmer, and Roman Vershynin. Private sampling: A noiseless approach for generating differentially private synthetic data. SIAM Journal on Mathematics of Data Science, 4(3):1082–1115, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1449944.

Barre:2022:CCV

[BTd22] Mathieu Barré, Adrien Taylor, and Alexandre d’Aspremont. Convergence of a constrained vector extrapolation scheme. SIAM Journal on Mathematics of Data Science, 4 (3):979–1002, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1428030.

Baldi:2019:PTF

[BV19] Pierre Baldi and Roman Vershynin. Polynomial threshold functions, hyperplane arrangements, and random tensors. SIAM Journal on Mathematics of Data Science, 1(4):699–729, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Calder:2019:CLL

[Cal19] Jeff Calder. Consistency of Lipschitz learning with infinite unlabeled data and finite labeled data. SIAM Journal on Mathematics of Data Science, 1(4): 780–812, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Charisopoulos:2021:CED

[CBD21] Vasileios Charisopoulos, Austin R. Benson, and Anil Damle. Communication-efficient distributed eigenspace estimation. SIAM Journal on Mathematics of Data Science, 3(4):1067– 1092, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Chambolle:2022:ABP

[CC22] Antonin Chambolle and Juan Pablo Contreras. Accelerated Bregman primal-dual methods applied to optimal transport and Wasserstein barycenter problems. SIAM Journal on Mathematics of Data Science, 4 (4):1369–1395, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1481865.

Carmona:2023:SOA

[CCNP23] Mikhael Carmona, Victor Chepoi, Guyslain Naves, and Pascal Préa. A simple and optimal algorithm for strict circular seriation. SIAM Journal on Mathematics of Data Science, 5(1):201–221, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1495342.

Cowen:2021:DSD

[CDH^{+}21] Lenore Cowen, Kapil Devkota, Xiaozhe Hu, James M. Murphy, and Kaiyi Wu. Diffusion state distances: Multitemporal analysis, fast algorithms, and applications to biological networks. SIAM Journal on Mathematics of Data Science, 3(1): 142–170, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Czaja:2021:TMF

[CDJN21] Wojciech Czaja, Dong Dong, Pierre-Emmanuel Jabin, and Franck Olivier Ndjakou Njeunje. Transport model for feature extraction. SIAM Journal on Mathematics of Data Science, 3(1):321–341, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Chen:2023:CSL

[CDS23] Wenyu Chen, Mathias Drton, and Ali Shojaie. Causal structural learning via local graphs. SIAM Journal on Mathematics of Data Science, 5(2):280–305, ???? 2023. CODEN SJMDAQ.

ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1362796.

Chodrow:2023:NSC

[CEH23] Philip Chodrow, Nicole Eikmeier, and Jamie Haddock. Nonbacktracking spectral clustering of nonuniform hypergraphs. SIAM Journal on Mathematics of Data Science, 5(2):251–279, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1494713.

Cohen:2019:ICD

[CG19] Jeremy E. Cohen and Nicolas Gillis. Identifiability of complete dictionary learning. SIAM Journal on Mathematics of Data Science, 1(3):518–536, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Chodrow:2020:MUR

[Cho20] Philip S. Chodrow. Moments of uniform random multigraphs with fixed degree sequences. SIAM Journal on Mathematics of Data Science, 2(4):1034– 1065, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Chevallier:2022:EWD

[CLLD22] Emmanuel Chevallier, Didong Li, Yulong Lu, and David Dunson. Exponential-wrapped distributions on symmetric spaces. SIAM Journal on Mathematics of Data Science, 4(4):1347–1368, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1461551.

Caragea:2022:QAR

[CLM^{+}22] Andrei Caragea, Dae Gwan Lee, Johannes Maly, Götz Pfander, and Felix Voigtlaender. Quantitative approximation results for complexvalued neural networks. SIAM Journal on Mathematics of Data Science, 4(2):553–580, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1429540.

Cuchiero:2020:DNN

[CLT20] Christa Cuchiero, Martin Larsson, and Josef Teichmann. Deep neural networks, generic universal interpolation, and controlled ODEs. SIAM Journal on Mathematics of Data Science, 2(3):901–919, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Chau:2021:SGL

[CMR^{+}21] Ngoc Huy Chau, Éric Moulines, Miklós Rásonyi, Sotirios Sabanis, and Ying Zhang. On stochastic gradient Langevin dynamics with dependent data streams: The fully nonconvex case. SIAM Journal on Mathematics of Data Science, 3(3): 959–986, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Clum:2020:MCA

[CMS20] Charles Clum, Dustin G.

Mixon, and Theresa Scarnati.

Matching component analysis for transfer learning. SIAM Journal on Mathematics of Data Science, 2(2):309–334, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Cooper:2021:GMO

[Coo21] Yaim Cooper. Global minima of overparameterized neural networks. SIAM Journal on Mathematics of Data Science, 3(2):676–691, ???? 2021. CO-DEN SJMDAQ. ISSN 2577-0187.

Combettes:2020:LCL

[CP20] Patrick L. Combettes and Jean-Christophe Pesquet. Lipschitz certificates for layered network structures driven by averaged activation operators. SIAM Journal on Mathematics of Data Science, 2(2):529–557, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Chen:2022:OOT

[CR22] Haolin Chen and Luis Rademacher. Overcomplete order-3 tensor decomposition, blind deconvolution, and Gaussian mixture models. SIAM Journal on Mathematics of Data Science, 4(1):336–361, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1399415.

Chi:2019:RTC

[CS19] Eric C. Chi and Stefan Steinerberger. Recovering trees with convex clustering. SIAM Journal on Mathematics of Data Science, 1(3):383–407, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Chen:2022:AEK

[CSAW22] Yuming Chen, Daniel Sanz-Alonso, and Rebecca Willett. Autodifferentiable ensemble Kalman filters. SIAM Journal on Mathematics of Data Science, 4(2):801–833, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1434477.

Cang:2020:PCD

[CW20] Zixuan Cang and Guo-Wei Wei. Persistent cohomology for data with multicomponent heterogeneous information. SIAM Journal on Mathematics of Data Science, 2(2): 396–418, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Cui:2020:ASN

[CZD^{+}20] Chunfeng Cui, Kaiqi Zhang, Talgat Daulbaev, Julia Gusak, Ivan Oseledets, and Zheng Zhang. Active subspace of neural networks: Structural analysis and universal attacks. SIAM Journal on Mathematics of Data Science, 2(4):1096– 1122, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Dar:2022:DDD

[DB22] Yehuda Dar and Richard G.

Baraniuk. Double double descent: On generalization errors in transfer learning between linear regression tasks. SIAM Journal on Mathematics of Data Science, 4(4):1447–1472, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1469559.

delBarrio:2023:ICL

[dBSLNW23] Eustasio del Barrio, Alberto González Sanz, Jean-Michel Loubes, and Jonathan Niles-Weed. An improved central limit theorem and fast convergence rates for entropic transportation costs. SIAM Journal on Mathematics of Data Science, 5(3):639–669, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M149260X.

DeBortoli:2021:MEM

[DDD^{+}21] Valentin De Bortoli, Agnès Desolneux, Alain Durmus, Bruno Galerne, and Arthur Leclaire. Maximum entropy methods for texture synthesis: Theory and practice. SIAM Journal on Mathematics of Data Science, 3(1):52–82, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

DeLoera:2020:STT

[DH20] Jesus A. De Loera and Thomas Hogan. Stochastic Tverberg theorems with applications in multiclass logistic regression, separability, and centerpoints of data. SIAM Journal on Mathematics of Data Science, 2(4):1151–1166, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Delgosha:2022:RCU

[DHP22] Payam Delgosha, Hamed Hassani, and Ramtin Pedarsani. Robust classification under ℓ_{0} attack for the Gaussian mixture model. SIAM Journal on Mathematics of Data Science, 4(1):362–385, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1426286.

Drees:2020:MDP

[DJRW20] Holger Drees, Anja Janßen, Sidney I. Resnick, and Tiandong Wang. On a minimum distance procedure for threshold selection in tail analysis. SIAM Journal on Mathematics of Data Science, 2(1):75–102, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Duruisseaux:2022:VFA

[DL22] Valentin Duruisseaux and Melvin Leok. A variational formulation of accelerated optimization on Riemannian manifolds. SIAM Journal on Mathematics of Data Science, 4(2):649–674, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1395648.

Dunlap:2022:LVS

[DM22] Alexander Dunlap and Jean-Christophe Mourrat. Local versions of sum-of-norms clustering. SIAM Journal on Mathematics of Data Science, 4(4):1250–1271, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1448732.

Doan:2021:FTP

[DMR21] Thinh T. Doan, Siva Theja Maguluri, and Justin Romberg. Finite-time performance of distributed temporal-difference learning with linear function approximation. SIAM Journal on Mathematics of Data Science, 3(1):298–320, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Dirksen:2022:SER

[DMS22] Sjoerd Dirksen, Shahar Mendelson, and Alexander Stollenwerk. Sharp estimates on random hyperplane tessellations. SIAM Journal on Mathematics of Data Science, 4 (4):1396–1419, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1485826.

Dietrich:2022:SDJ

[DYM^{+}22] Felix Dietrich, Or Yair, Rotem Mulayoff, Ronen Talmon, and Ioannis G. Kevrekidis. Spectral discovery of jointly smooth features for multimodal data. SIAM Journal on Mathematics of Data Science, 4(1):410–430, ???? 2022. CODEN SJMDAQ.

ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M141590X.

Ghaoui:2021:IDL

[EGT^{+}21] Laurent El Ghaoui, Fangda Gu, Bertrand Travacca, Armin Askari, and Alicia Tsai. Implicit deep learning. SIAM Journal on Mathematics of Data Science, 3(3):930–958, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Elvira:2023:SRI

[EH23] Clément Elvira and Cédric Herzet. Safe rules for the identification of zeros in the solutions of the SLOPE problem. SIAM Journal on Mathematics of Data Science, 5(1):147–173, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1457631.

ElAlaoui:2019:DPD

[ERK^{+}19] Ahmed El Alaoui, Aaditya Ramdas, Florent Krzakala, Lenka Zdeborová, and Michael I. Jordan. Decoding from pooled data: Sharp informationtheoretic bounds. SIAM Journal on Mathematics of Data Science, 1(1):161–188, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Etter:2022:OSG

[EY22] Philip A. Etter and Lexing Ying. Operator shifting for general noisy matrix systems. SIAM Journal on Mathematics of Data Science, 4(4):1320–1346, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1416849.

Fanuel:2022:PSD

[FADS22] Michaël Fanuel, Antoine Aspeel, Jean-Charles Delvenne, and Johan A. K. Suykens. Positive semi-definite embedding for dimensionality reduction and out-of-sample extensions. SIAM Journal on Mathematics of Data Science, 4(1):153–178, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/20M1370653.

Fefferman:2020:RRM

[FILN20] Charles Fefferman, Sergei Ivanov, Matti Lassas, and Hariharan Narayanan. Reconstruction of a Riemannian manifold from noisy intrinsic distances. SIAM Journal on Mathematics of Data Science, 2(3):770–808, ???? 2020. CO-DEN SJMDAQ. ISSN 2577-0187.

Fan:2022:JCD

[FKZ22] Yifeng Fan, Yuehaw Khoo, and Zhizhen Zhao. Joint community detection and rotational synchronization via semidefinite programming. SIAM Journal on Mathematics of Data Science, 4(3):1052–1081, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1419702.

Fanuel:2021:DSI

[FSS21] Michael Fanuel, Joachim Schreurs, and Johan Suykens. Diversity sampling is an implicit regularization for kernel methods. SIAM Journal on Mathematics of Data Science, 3(1):280–297, ???? 2021. CO-DEN SJMDAQ. ISSN 2577-0187.

Fanuel:2022:DPP

[FSS22] Michaël Fanuel, Joachim Schreurs, and Johan A. K. Suykens. Determinantal point processes implicitly regularize semiparametric regression problems. SIAM Journal on Mathematics of Data Science, 4(3):1171–1190, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1403977.

Fasino:2020:ECH

[FT20] Dario Fasino and Francesco Tudisco. Ergodicity coefficients for higher-order stochastic processes. SIAM Journal on Mathematics of Data Science, 2(3): 740–769, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Farnia:2023:GGG

[FWDJ23] Farzan Farnia, William W.

Wang, Subhro Das, and Ali Jadbabaie. GAT-GMM: Generative adversarial training for Gaussian mixture models. SIAM Journal on Mathematics of Data Science, 5(1):122–146, ???? 2023. CODEN SJMDAQ.

ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1445831.

Gupta:2020:SLP

[GCPT20] Abhishek Gupta, Hao Chen, Jianzong Pi, and Gaurav Tendolkar. Some limit properties of Markov chains induced by recursive stochastic algorithms. SIAM Journal on Mathematics of Data Science, 2(4):967–1003, ???? 2020. CO-DEN SJMDAQ. ISSN 2577-0187.

Gu:2021:MFC

[GGWX21] Haotian Gu, Xin Guo, Xiaoli Wei, and Renyuan Xu. Meanfield controls with Q-learning for cooperative MARL: Convergence and complexity analysis. SIAM Journal on Mathematics of Data Science, 3(4): 1168–1196, ???? 2021. CO-DEN SJMDAQ. ISSN 2577-0187.

Gupta:2021:CRS

[GH21] Abhishek Gupta and William B. Haskell. Convergence of recursive stochastic algorithms using Wasserstein divergence. SIAM Journal on Mathematics of Data Science, 3(4):1141– 1167, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Gidisu:2022:GCD

[GH22] Perfect Y. Gidisu and Michiel E. Hochstenbach. A generalized CUR decomposition for matrix pairs. SIAM Journal on Mathematics of Data Science, 4(1):386–409, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1432119.

Grishchenko:2021:DLS

[GIMA21] Dmitry Grishchenko, Franck Iutzeler, Jérôme Malick, and Massih-Reza Amini. Distributed learning with sparse communications by identification. SIAM Journal on Mathematics of Data Science, 3(2): 715–735, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Gao:2019:GPLb

[GKBD19] Tingran Gao, Shahar Z. Kovalsky, Doug M. Boyer, and Ingrid Daubechies. Gaussian process landmarking for threedimensional geometric morphometrics. SIAM Journal on Mathematics of Data Science, 1(1):237–267, ???? 2019. CO-DEN SJMDAQ. ISSN 2577-0187.

Gao:2019:GPLa

[GKD19] Tingran Gao, Shahar Z. Kovalsky, and Ingrid Daubechies. Gaussian process landmarking on manifolds. SIAM Journal on Mathematics of Data Science, 1(1):208–236, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Gossmann:2021:TDR

[GPWS21] Alexej Gossmann, Aria Pezeshk, Yu-Ping Wang, and Berkman Sahiner. Test data reuse for the evaluation of continuously evolving classification algorithms using the area under the receiver operating characteristic curve. SIAM Journal on Mathematics of Data Science, 3(2):692–714, ???? 2021. CO-DEN SJMDAQ. ISSN 2577-0187.

Gunther:2020:LPT

[GRS^{+}20] Stefanie Günther, Lars Ruthotto, Jacob B. Schroder, Eric C. Cyr, and Nicolas R. Gauger. Layer-parallel training of deep residual neural networks. SIAM Journal on Mathematics of Data Science, 2(1):1–23, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Heaton:2022:WBP

[HFL^{+}22] Howard Heaton, Samy Wu Fung, Alex Tong Lin, Stanley Osher, and Wotao Yin. Wasserstein-based projections with applications to inverse problems. SIAM Journal on Mathematics of Data Science, 4(2):581–603, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/20M1376790.

Habring:2022:GVM

[HH22] Andreas Habring and Martin Holler. A generative variational model for inverse problems in imaging. SIAM Journal on Mathematics of Data Science, 4(1):306–335, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1414978.

Habring:2023:NRI

[HH23] Andreas Habring and Martin Holler. A note on the regularity of images generated by convolutional neural networks. SIAM Journal on Mathematics of Data Science, 5(3):670–692, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1525995.

Hamm:2023:WWI

[HHK23] Keaton Hamm, Nick Henscheid, and Shujie Kang. Wassmap: Wasserstein isometric mapping for image manifold learning. SIAM Journal on Mathematics of Data Science, 5(2):475–501, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1490053.

Huang:2022:DRR

[HHV22] Ningyuan Teresa Huang, David W.

Hogg, and Soledad Villar. Dimensionality reduction, regularization, and generalization in overparameterized regressions. SIAM Journal on Mathematics of Data Science, 4(1):126–152, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/20M1387821.

Hanik:2022:BID

[HHvT22] Martin Hanik, Hans-Christian Hege, and Christoph von Tycowicz. Bi-invariant dissimilarity measures for sample distributions in Lie groups. SIAM Journal on Mathematics of Data Science, 4(4):1223–1249, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1410373.

Horvath:2022:ASG

[HLRJ22] Samuel Horváth, Lihua Lei, Peter Richtárik, and Michael I. Jordan. Adaptivity of stochastic gradient methods for nonconvex optimization. SIAM Journal on Mathematics of Data Science, 4(2):634–648, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1394308.

Hou:2020:AAE

[HLZ20] Thomas Y. Hou, Zhenzhen Li, and Ziyun Zhang. Analysis of asymptotic escape of strict saddle sets in manifold optimization. SIAM Journal on Mathematics of Data Science, 2(3): 840–871, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Hawkins:2022:TCN

[HLZ22] Cole Hawkins, Xing Liu, and Zheng Zhang. Towards compact neural networks via endto-end training: a Bayesian tensor approach with automatic rank determination.

SIAM Journal on Mathematics of Data Science, 4(1):46–71, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1391444.

Haddock:2021:GWI

[HM21] Jamie Haddock and Anna Ma. Greed works: an improved analysis of sampling Kaczmarz–Motzkin. SIAM Journal on Mathematics of Data Science, 3(1):342–368, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Heinemann:2022:RWB

[HMZ22] Florian Heinemann, Axel Munk, and Yoav Zemel. Randomized Wasserstein barycenter computation: Resampling with statistical guarantees. SIAM Journal on Mathematics of Data Science, 4(1):229–259, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/20M1385263.

Ho:2019:SSI

[HN19] Nhat Ho and XuanLong Nguyen. Singularity structures and impacts on parameter estimation in finite mixtures of distributions. SIAM Journal on Mathematics of Data Science, 1(4):730–758, ???? 2019. CO-DEN SJMDAQ. ISSN 2577-0187.

Hickok:2022:ASS

[HNP22] Abigail Hickok, Deanna Needell, and Mason A. Porter. Analysis of spatial and spatiotemporal anomalies using persistent homology: Case studies with COVID-19 data. SIAM Journal on Mathematics of Data Science, 4(3):1116–1144, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1435033.

Hien:2022:BBM

[HPG^{+}22] Le Thi Khanh Hien, Duy Nhat Phan, Nicolas Gillis, Masoud Ahookhosh, and Panagiotis Patrinos. Block Bregman majorization minimization with extrapolation. SIAM Journal on Mathematics of Data Science, 4(1):1–25, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1432661.

Hamm:2022:IDA

[HS22] Thomas Hamm and Ingo Steinwart. Intrinsic dimension adaptive partitioning for kernel methods. SIAM Journal on Mathematics of Data Science, 4(2):721–749, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1435690.

Huguet:2023:TID

[HTR^{+}23] Guillaume Huguet, Alexander Tong, Bastian Rieck, Jessie Huang, Manik Kuchroo, Matthew Hirn, Guy Wolf, and Smita Krishnaswamy. Timeinhomogeneous diffusion geometry and topology. SIAM Journal on Mathematics of Data Science, 5(2):346–372, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1462945.

Hong:2023:OWP

[HYFB23] David Hong, Fan Yang, Jeffrey A. Fessler, and Laura Balzano. Optimally weighted PCA for high-dimensional heteroscedastic data. SIAM Journal on Mathematics of Data Science, 5(1):222–250, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1470244.

Harris:2021:DTC

[HZ21] Kameron Decker Harris and Yizhe Zhu. Deterministic tensor completion with hypergraph expanders. SIAM Journal on Mathematics of Data Science, 3(4):1117–1140, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Jaffe:2021:SNJ

[JAA^{+}21] Ariel Jaffe, Noah Amsel, Yariv Aizenbud, Boaz Nadler, Joseph T. Chang, and Yuval Kluger. Spectral neighbor joining for reconstruction of latent tree models. SIAM Journal on Mathematics of Data Science, 3(1):113–141, ???? 2021. CO-DEN SJMDAQ. ISSN 2577-

Jia:2022:UGM

[JB22] Junteng Jia and Austin R. Benson. A unifying generative model for graph learning algorithms: Label propagation, graph convolutions, and combinations. SIAM Journal on Mathematics of Data Science, 4(1):100–125, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1395351.

Jiang:2023:OCN

[JC23] Jie Jiang and Xiaojun Chen. Optimality conditions for nonsmooth nonconvex-nonconcave min-max problems and generative adversarial networks. SIAM Journal on Mathematics of Data Science, 5(3):693–722, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1482238.

Jiang:2023:ARM

[JCD23] Liwei Jiang, Yudong Chen, and Lijun Ding. Algorithmic regularization in modelfree overparametrized asymmetric matrix factorization. SIAM Journal on Mathematics of Data Science, 5(3):723–744, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1519833.

Jorgensen:2021:APR

[JS21] Palle Jorgensen and David E. Stewart. Approximation properties of ridge functions and extreme learning machines. SIAM Journal on Mathematics of Data Science, 3(3):815–832, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Kolda:2020:SGL

[KH20] Tamara G. Kolda and David Hong. Stochastic gradients for large-scale tensor decomposition. SIAM Journal on Mathematics of Data Science, 2(4): 1066–1095, ???? 2020. CO-DEN SJMDAQ. ISSN 2577-0187.

Kolda:2019:ISJ

[Kol19] Tamara G. Kolda. Introduction to SIAM Journal on Mathematics of Data Science (SIMODS). SIAM Journal on Mathematics of Data Science, 1(1):1–7, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Khamaru:2021:TDL

[KPR^{+}21] Koulik Khamaru, Ashwin Pananjady, Feng Ruan, Martin J. Wainwright, and Michael I. Jordan. Is temporal difference learning optimal? an instancedependent analysis. SIAM Journal on Mathematics of Data Science, 3(4):1013–1040, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Klebanov:2020:RTC

[KSS20] Ilja Klebanov, Ingmar Schuster, and T. J. Sullivan. A rigorous theory of conditional mean embeddings. SIAM Journal on Mathematics of Data Science, 2(3):583–606, ???? 2020. CO-DEN SJMDAQ. ISSN 2577-0187.

Kueng:2021:BCD

[KT21] Richard Kueng and Joel A. Tropp. Binary component decomposition. Part I: The positive-semidefinite case. SIAM Journal on Mathematics of Data Science, 3(2):544–572, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Klatt:2020:ERO

[KTM20] Marcel Klatt, Carla Tameling, and Axel Munk. Empirical regularized optimal transport: Statistical theory and applications. SIAM Journal on Mathematics of Data Science, 2(2): 419–443, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Kuo:2020:GSS

[KZLW20] Han-Wen Kuo, Yuqian Zhang, Yenson Lau, and John Wright. Geometry and symmetry in short-and-sparse deconvolution. SIAM Journal on Mathematics of Data Science, 2(1): 216–245, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Landa:2023:RIM

[LC23] Boris Landa and Xiuyuan Cheng. Robust inference of manifold density and geometry by doubly stochastic scaling. SIAM Journal on Mathematics of Data Science, 5(3):589–614, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1516968.

Landa:2021:DSN

[LCK21] Boris Landa, Ronald R. Coifman, and Yuval Kluger. Doubly stochastic normalization of the Gaussian kernel is robust to heteroskedastic noise. SIAM Journal on Mathematics of Data Science, 3(1):388–413, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Leeb:2021:MDW

[Lee21] William E. Leeb. Matrix denoising for weighted loss functions and heterogeneous signals. SIAM Journal on Mathematics of Data Science, 3(3): 987–1012, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Li:2021:GEM

[Li21]  Weilin Li.  Generalization

error of minimum weighted norm and kernel interpolation. SIAM Journal on Mathematics of Data Science, 3(1):414–438, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Li:2021:MGM

[LKA^{+}21] Qunwei Li, Bhavya Kailkhura, Rushil Anirudh, Jize Zhang, Yi Zhou, Yingbin Liang, T. Yong-Jin Han, and Pramod K. Varshney. MR-AN: Manifold regularized generative adversarial networks for scientific data. SIAM Journal on Mathematics of Data Science, 3(4): 1197–1222, ???? 2021. CO-DEN SJMDAQ. ISSN 2577-0187.

Li:2022:DCO

[LLC22] Boyue Li, Zhize Li, and Yuejie Chi. DESTRESS: Computation-optimal and communicationefficient decentralized nonconvex finite-sum optimization. SIAM Journal on Mathematics of Data Science, 4 (3):1031–1051, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1450677.

Li:2022:NWD

[LLM22] Yuqing Li, Tao Luo, and Chao Ma. Nonlinear weighted directed acyclic graph and a priori estimates for neural networks. SIAM Journal on Mathematics of Data Science, 4(2):694–720, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M140955X.

Li:2023:AIO

[LLP23] Tianjiao Li, Guanghui Lan, and Ashwin Pananjady. Accelerated and instance-optimal policy evaluation with linear function approximation. SIAM Journal on Mathematics of Data Science, 5(1):174–200, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1468668.

Lovas:2023:TNN

[LLRS23] Attila Lovas, Iosif Lytras, Miklós Rásonyi, and Sotirios Sabanis. Taming neural networks with TUSLA: Nonconvex learning via adaptive stochastic gradient Langevin algorithms. SIAM Journal on Mathematics of Data Science, 5(2):323–345, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1514283.

Lai:2020:CSC

[LM20] Ming-Jun Lai and Daniel Mckenzie. Compressive sensing for cut improvement and local clustering. SIAM Journal on Mathematics of Data Science, 2(2):368–395, ???? 2020. CO-DEN SJMDAQ. ISSN 2577-0187.

Little:2022:BGD

[LMM22] Anna Little, Daniel McKenzie, and James M. Murphy. Balancing geometry and density: Path distances on highdimensional data. SIAM Journal on Mathematics of Data Science, 4(1):72–99, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/20M1386657.

Li:2020:GBR

[LMR^{+}20] Yuan Li, Benjamin Mark, Garvesh Raskutti, Rebecca Willett, Hyebin Song, and David Neiman. Graphbased regularization for regression problems with alignment and highly correlated designs. SIAM Journal on Mathematics of Data Science, 2(2):480–504, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Luo:2022:ECL

[LMXZ22] Tao Luo, Zheng Ma, Zhi-Qin John Xu, and Yaoyu Zhang. On the exact computation of linear frequency principle dynamics and its generalization. SIAM Journal on Mathematics of Data Science, 4(4):1272–1292, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1444400.

Li:2023:MOR

[LPH^{+}23] Adam Li, Ronan Perry, Chester Huynh, Tyler M. Tomita, Ronak Mehta, Jesus Arroyo, Jesse Patsolic, Ben Falk, Sridevi Sarma, and Joshua Vogelstein. Manifold oblique random forests: Towards closing the gap on convolutional deep networks. SIAM Journal on Mathematics of Data Science, 5(1):77–96, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1449117.

Lee:2019:LFD

[LR19] Jaeho Lee and Maxim Raginsky. Learning finitedimensional coding schemes with nonlinear reconstruction maps. SIAM Journal on Mathematics of Data Science, 1(3): 617–642, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Linderman:2019:CSP

[LS19] George C. Linderman and Stefan Steinerberger. Clustering with t-SNE, provably. SIAM Journal on Mathematics of Data Science, 1(2):313–332, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Lee:2023:RIA

[LS23] Kiryung Lee and Dominik Stöger. Randomly initialized alternating least squares: Fast convergence for matrix sensing. SIAM Journal on Mathematics of Data Science, 5(3):774–799, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1506456.

Luke:2019:OSM

[LST19] D. Russell Luke, Shoham Sabach, and Marc Teboulle. Optimization on spheres: Models and proximal algorithms with computational performance comparisons. SIAM Journal on Mathematics of Data Science, 1(3):408–445, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Lopes:2020:MAC

[LWL20] Miles E. Lopes, Suofei Wu, and Thomas C. M. Lee. Measuring the algorithmic convergence of randomized ensembles: The regression setting. SIAM Journal on Mathematics of Data Science, 2(4):921–943, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Landa:2022:BRR

[LZK22] Boris Landa, Thomas T. C. K. Zhang, and Yuval Kluger. Biwhitening reveals the rank of a count matrix. SIAM Journal on Mathematics of Data Science, 4(4):1420–1446, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1456807.

Montanelli:2019:NEB

[MD19] Hadrien Montanelli and Qiang Du. New error bounds for deep ReLU networks using sparse grids. SIAM Journal on Mathematics of Data Science, 1(1): 78–92, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Molina-Fructuoso:2022:TDH

[MFM22] Martin Molina-Fructuoso and Ryan Murray. Tukey depths and Hamilton–Jacobi differential equations. SIAM Journal on Mathematics of Data Science, 4(2):604–633, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1411998.

Marchand:2021:ANP

[MGHV21] Melissa Marchand, Kyle Gallivan, Wen Huang, and Paul Van Dooren. Analysis of the neighborhood pattern similarity measure for the role extraction problem. SIAM Journal on Mathematics of Data Science, 3(2):736–757, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Malgouyres:2019:MCS

[ML19] François Malgouyres and Joseph Landsberg. Multilinear compressive sensing and an application to convolutional linear networks. SIAM Journal on Mathematics of Data Science, 1(3):446–475, ???? 2019. CO-DEN SJMDAQ. ISSN 2577-0187.

Mordant:2023:SAR

[MM23] Gilles Mordant and Axel Munk. Statistical analysis of random objects via metric measure Laplacians. SIAM Journal on Mathematics of Data Science, 5(2):528–557, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1491022.

Maroulas:2020:BFP

[MNO20] Vasileios Maroulas, Farzana Nasrin, and Christopher Oballe. A Bayesian framework for persistent homology. SIAM Journal on Mathematics of Data Science, 2(1):48–74, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Mukkamala:2020:CCB

[MOPS20] Mahesh Chandra Mukkamala, Peter Ochs, Thomas Pock, and Shoham Sabach. Convexconcave backtracking for inertial Bregman proximal gradient algorithms in nonconvex optimization. SIAM Journal on Mathematics of Data Science, 2(3):658–682, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Mazumdar:2020:GBL

[MRS20] Eric Mazumdar, Lillian J. Ratliff, and S. Shankar Sastry. On gradient-based learning in continuous games. SIAM Journal on Mathematics of Data Science, 2(1):103–131, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Minster:2020:RAL

[MSK20] Rachel Minster, Arvind K. Saibaba, and Misha E. Kilmer. Randomized algorithms for low-rank tensor decompositions in the Tucker format. SIAM Journal on Mathematics of Data Science, 2(1):189–215, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Maire:2022:MKL

[MV22] Florian Maire and Pierre Vandekerkhove. Markov kernels local aggregation for noise vanishing distribution sampling. SIAM Journal on Mathematics of Data Science, 4 (4):1293–1319, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1469626.

Memoli:2022:PLP

[MWW22] Facundo Mémoli, Zhengchao Wan, and Yusu Wang. Persistent Laplacians: Properties, algorithms and implications. SIAM Journal on Mathematics of Data Science, 4(2):858–884, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1435471.

Neumayer:2023:ALF

[NGBU23] Sebastian Neumayer, Alexis Goujon, Pakshal Bohra, and Michael Unser. Approximation of Lipschitz functions using deep spline neural networks. SIAM Journal on Mathematics of Data Science, 5(2):306–322, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1504573.

Newman:2021:TLV

[NRHvBW21] Elizabeth Newman, Lars Ruthotto, Joseph Hart, and Bart van Bloemen Waanders. Train like a (var)Pro: Efficient training of neural networks with variable projection. SIAM Journal on Mathematics of Data Science, 3(4):1041– 1066, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Ongie:2021:TMN

[OPAB^{+}21] Greg Ongie, Daniel Pimentel-Alarcón, Laura Balzano, Rebecca Willett, and Robert D. Nowak. Tensor methods for nonlinear matrix completion.

SIAM Journal on Mathematics of Data Science, 3(1):253–279, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

OReilly:2022:SGG

[OT22] Eliza O’Reilly and Ngoc Mai Tran. Stochastic geometry to generalize the Mondrian process. SIAM Journal on Mathematics of Data Science, 4(2):531–552, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1354490.

Osting:2021:CAA

[OWXZ21] Braxton Osting, Dong Wang, Yiming Xu, and Dominique Zosso. Consistency of archetypal analysis. SIAM Journal on Mathematics of Data Science, 3(1):1–30, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Pamfil:2019:RMM

[PHLP19] A. Roxana Pamfil, Sam D. Howison, Renaud Lambiotte, and Mason A. Porter. Relating modularity maximization and stochastic block models in multilayer networks. SIAM Journal on Mathematics of Data Science, 1(4):667–698, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Parhi:2022:WKF

[PN22] Rahul Parhi and Robert D. Nowak. What kinds of functions do deep neural networks learn? insights from variational spline theory. SIAM Journal on Mathematics of Data Science, 4(2):464–489, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1418642.

Potts:2021:IAH

[PS21] Daniel Potts and Michael Schmischke. Interpretable approximation of high-dimensional data. SIAM Journal on Mathematics of Data Science, 3(4): 1301–1323, ???? 2021. CO-DEN SJMDAQ. ISSN 2577-0187.

Perry:2019:SCM

[PWB^{+}19] Amelia Perry, Jonathan Weed, Afonso S. Bandeira, Philippe Rigollet, and Amit Singer. The sample complexity of multireference alignment. SIAM Journal on Mathematics of Data Science, 1(3):497–517, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Romanov:2021:MRA

[RBO21] Elad Romanov, Tamir Bendory, and Or Ordentlich. Multi-reference alignment in high dimensions: Sample complexity and phase transition. SIAM Journal on Mathematics of Data Science, 3(2):494–523, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Reeves:2020:NI

[RDB20] Thomas Reeves, Anil Damle, and Austin R. Benson. Network interpolation. SIAM Journal on Mathematics of Data Science, 2(2):505–528, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Riihimaki:2023:SCD

[Rii23]  Henri Riihimäki.  Simpli-

cial q-connectivity of directed graphs with applications to network analysis. SIAM Journal on Mathematics of Data Science, 5(3):800–828, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1480021.

Roald:2022:AAA

[RSC^{+}22] Marie Roald, Carla Schenker, Vince D. Calhoun, Tülay Adali, Rasmus Bro, Jeremy E. Cohen, and Evrim Acar. An AO-ADMM approach to constraining PARAFAC2 on all modes. SIAM Journal on Mathematics of Data Science, 4(3):1191–1222, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1450033.

Saul:2022:NMD

[Sau22] Lawrence K. Saul. A nonlinear matrix decomposition for mining the zeros of sparse data. SIAM Journal on Mathematics of Data Science, 4(2):431–463, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1405769.

Shnitzer:2019:RHC

[SBCG^{+}19] Tal Shnitzer, Mirela Ben-Chen, Leonidas Guibas, Ronen Talmon, and Hau-Tieng Wu. Recovering hidden components in multimodal data with composite diffusion operators. SIAM Journal on Mathematics of Data Science, 1(3):588–616, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Salvi:2021:SKS

[SCF^{+}21] Cristopher Salvi, Thomas Cass, James Foster, Terry Lyons, and Weixin Yang. The signature kernel is the solution of a Goursat PDE. SIAM Journal on Mathematics of Data Science, 3(3):873–899, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Sun:2020:LRT

[SGL^{+}20] Yiming Sun, Yang Guo, Charlene Luo, Joel Tropp, and Madeleine Udell. Low-rank Tucker approximation of a tensor from streaming data. SIAM Journal on Mathematics of Data Science, 2(4):1123–1150, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Solomon:2022:VCN

[SGN22] Justin Solomon, Kristjan Greenewald, and Haikady Nagaraja. k-variance: A clustered notion of variance. SIAM Journal on Mathematics of Data Science, 4(3):957–978, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1385895.

Schaeffer:2020:ESS

[SM20] Hayden Schaeffer and Scott G. McCalla. Extending the stepsize restriction for gradient descent to avoid strict saddle points. SIAM Journal on Mathematics of Data Science, 2(4):1181–1197, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Shinde:2021:MES

[SNS21] Nimita Shinde, Vishnu Narayanan, and James Saunderson. Memoryefficient structured convex optimization via extreme point sampling. SIAM Journal on Mathematics of Data Science, 3(3):787–814, ???? 2021. CO-DEN SJMDAQ. ISSN 2577-0187.

Shin:2021:BRC

[SRR21] Jaehyeok Shin, Aaditya Ramdas, and Alessandro Rinaldo. On the bias, risk, and consistency of sample means in multi-armed bandits. SIAM Journal on Mathematics of Data Science, 3(4):1278–1300, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Schaub:2020:BIS

[SST20] Michael T. Schaub, Santiago Segarra, and John N. Tsitsiklis. Blind identification of stochastic block models from dynamical observations. SIAM Journal on Mathematics of Data Science, 2(2):335–367, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Sprangers:2023:GIT

[SV23] Brent Sprangers and Nick Vannieuwenhoven. Groupinvariant tensor train networks for supervised learning. SIAM Journal on Mathematics of Data Science, 5(4):829–853, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1506857.

Torres:2021:NEU

[TCTER21] Leo Torres, Kevin S. Chan, Hanghang Tong, and Tina Eliassi-Rad. Nonbacktracking eigenvalues under node removal: X-centrality and targeted immunization. SIAM Journal on Mathematics of Data Science, 3(2):656–675, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Tudisco:2019:NSM

[TH19] Francesco Tudisco and Desmond J.

Higham. A nonlinear spectral method for Core–Periphery detection in networks. SIAM Journal on Mathematics of Data Science, 1(2):269–292, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Tudisco:2023:CPD

[TH23] Francesco Tudisco and Desmond J.

Higham. Core-periphery detection in hypergraphs. SIAM Journal on Mathematics of Data Science, 5(1):1–21, ???? 2023. CODEN SJMDAQ.

ISSN 2577-0187. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1480926.

Tank:2021:CMD

[TLFS21] Alex Tank, Xiudi Li, Emily B. Fox, and Ali Shojaie. The convex mixture distribution: Granger causality for categorical time series. SIAM Journal on Mathematics of Data Science, 3(1):83–112, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Thao:2020:PRS

[TLSV20] Nguyen Hieu Thao, David Russell Luke, Oleg Soloviev, and Michel Verhaegen. Phase retrieval with sparse phase constraint. SIAM Journal on Mathematics of Data Science, 2(1):246–263, ???? 2020. CO-DEN SJMDAQ. ISSN 2577-0187.

Trillos:2020:MPA

[TM20] Nicolas García Trillos and Ryan W. Murray. A maximum principle argument for the uniform convergence of graph Laplacian regressors. SIAM Journal on Mathematics of Data Science, 2(3):705–739, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Tan:2023:DDM

[TMTS23] Hong Ye Tan, Subhadip Mukherjee, Junqi Tang, and Carola-Bibiane Schönlieb. Datadriven mirror descent with input-convex neural networks. SIAM Journal on Mathematics of Data Science, 5(2):558–587, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1508613.

Trillos:2019:VLN

[Tri19] Nicolas Garcia Trillos. Variational limits of k-NN graphbased functionals on data clouds. SIAM Journal on Mathematics of Data Science, 1(1):93–120, ???? 2019. CO-DEN SJMDAQ. ISSN 2577-0187.

Tanner:2019:MRI

[TTV19] Jared Tanner, Andrew Thompson, and Simon Vary. Matrix rigidity and the ill-posedness of robust PCA and matrix completion. SIAM Journal on Mathematics of Data Science, 1(3):537–554, ???? 2019. CO-DEN SJMDAQ. ISSN 2577-0187.

Udell:2019:WBD

[UT19] Madeleine Udell and Alex Townsend. Why are big data matrices approximately low rank? SIAM Journal on Mathematics of Data Science, 1(1): 144–160, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Vershynin:2020:MCN

[Ver20] Roman Vershynin. Memory capacity of neural networks with threshold and rectified linear unit activations. SIAM Journal on Mathematics of Data Science, 2(4):1004–1033, ????

2020. CODEN SJMDAQ. ISSN 2577-0187.

Valdeira:2023:PRG

[VFMS23] Filipa M. Valdeira, Ricardo Ferreira, Alessandra Micheletti, and Cláudia Soares. Probabilistic registration for Gaussian process three-dimensional shape modelling in the presence of extensive missing data. SIAM Journal on Mathematics of Data Science, 5(2):502–527, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1495494.

Veldt:2019:MCO

[VGWS19] Nate Veldt, David F. Gleich, Anthony Wirth, and James Saunderson. Metricconstrained optimization for graph clustering algorithms. SIAM Journal on Mathematics of Data Science, 1(2):333–355, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Wang:2022:BCG

[WT22] Ke Wang and Christos Thrampoulidis. Binary classification of Gaussian mixtures: Abundance of support vectors, benign overfitting, and regularization. SIAM Journal on Mathematics of Data Science, 4(1):260–284, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1415121.

Webber:2021:EBD

[WTD^{+}21] Robert J. Webber, Erik H.

Thiede, Douglas Dow, Aaron R. Dinner, and Jonathan Weare. Error bounds for dynamical spectral estimation. SIAM Journal on Mathematics of Data Science, 3(1):225–252, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Weir:2020:MMB

[WWZM20] William H. Weir, Benjamin Walker, Lenka Zdeborová, and Peter J. Mucha. Multilayer modularity belief propagation to assess detectability of community structure. SIAM Journal on Mathematics of Data Science, 2(3):872–900, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Wang:2020:ERE

[WYSO20] Bao Wang, Binjie Yuan, Zuoqiang Shi, and Stanley J. Osher. EnResNet: ResNets ensemble via the Feynman–Kac formalism for adversarial defense and beyond. SIAM Journal on Mathematics of Data Science, 2(3):559–582, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Xie:2022:OGE

[XCRW22] Yuege Xie, Hung-Hsu Chou, Holger Rauhut, and Rachel Ward. Overparameterization and generalization error: Weighted trigonometric interpolation. SIAM Journal on Mathematics of Data Science, 4(2):885–908, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1390955.

Xing:2022:BIN

[XSC22] Yue Xing, Qifan Song, and Guang Cheng. Benefit of interpolation in nearest neighbor algorithms. SIAM Journal on Mathematics of Data Science, 4(2):935–956, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1437457.

Yongacoglu:2023:SPI

[YAY23] Bora Yongacoglu, Gürdal Arslan, and Serdar Yüksel. Satisficing paths and independent multiagent reinforcement learning in stochastic games. SIAM Journal on Mathematics of Data Science, 5(3):745–773, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1515112.

Yuan:2019:MSH

[YLB^{+}19] Baichuan Yuan, Hao Li, Andrea L. Bertozzi, P. Jeffrey Brantingham, and Mason A. Porter. Multivariate spatiotemporal Hawkes processes and network reconstruction. SIAM Journal on Mathematics of Data Science, 1(2):356–382, ???? 2019. CODEN SJMDAQ. ISSN 2577-0187.

Yurtsever:2021:SSP

[YTF^{+}21] Alp Yurtsever, Joel A. Tropp, Olivier Fercoq, Madeleine Udell, and Volkan Cevher. Scalable semidefinite programming. SIAM Journal on Mathematics of Data Science, 3(1): 171–200, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Zhang:2021:NBB

[ZG21] Ruda Zhang and Roger Ghanem.

Normal-bundle bootstrap. SIAM Journal on Mathematics of Data Science, 3(2):573–592, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.

Zhang:2022:FGT

[Zha22]  Tong Zhang.  Feel-good

Thompson sampling for contextual bandits and reinforcement learning. SIAM Journal on Mathematics of Data Science, 4(2):834–857, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M140924X.

Zhang:2020:IFO

[ZLRY20] Anru R. Zhang, Yuetian Luo, Garvesh Raskutti, and Ming Yuan. ISLET: Fast and optimal low-rank tensor regression via importance sketching. SIAM Journal on Mathematics of Data Science, 2(2):444–479, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Zhang:2020:DOC

[ZLZ20] Yuan Zhang, Elizaveta Levina, and Ji Zhu. Detecting overlapping communities in networks using spectral methods. SIAM Journal on Mathematics of Data Science, 2(2):265–283, ???? 2020. CODEN SJMDAQ. ISSN 2577-0187.

Zilber:2022:GPO

[ZN22] Pini Zilber and Boaz Nadler. GNMR: A provable one-line algorithm for low rank matrix recovery. SIAM Journal on Mathematics of Data Science, 4(2):909–934, ???? 2022. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/21M1433812.

Zhu:2019:SMD

[ZQW^{+}19] Wei Zhu, Qiang Qiu, Bao Wang, Jianfeng Lu, Guillermo Sapiro, and Ingrid Daubechies. Stop memorizing: a datadependent regularization framework for intrinsic pattern learning. SIAM Journal on Mathematics of Data Science, 1(3):476–496, ???? 2019. CO-DEN SJMDAQ. ISSN 2577-0187.

Zheng:2023:EIB

[ZRG23] Léon Zheng, Elisa Riccietti, and Rémi Gribonval. Efficient identification of butterfly sparse matrix factorizations. SIAM Journal on Mathematics of Data Science, 5(1):22–49, ???? 2023. CODEN SJMDAQ.

ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1488727.

Zheng:2020:SSL

[ZYWL20] Xinye Zheng, Jianbo Ye, James Z. Wang, and Jia Li. SCOTT: Shape-location combined tracking with optimal transport. SIAM Journal on Mathematics of Data Science, 2(2):284–308, ???? 2020. CO-DEN SJMDAQ. ISSN 2577-0187.

Zhang:2023:PTQ

[ZZS23] Jinjie Zhang, Yixuan Zhou, and Rayan Saab. Post-training quantization for neural networks with provable guarantees. SIAM Journal on Mathematics of Data Science, 5(2):373–399, ???? 2023. CODEN SJMDAQ. ISSN 2577-0187. URL https:

//epubs.siam.org/doi/10.
1137/22M1511709.

Zhang:2021:ELC

[ZZST21] Ning Zhang, Yangjing Zhang, Defeng Sun, and Kim-Chuan Toh. An efficient linearly convergent regularized proximal point algorithm for fused multiple graphical lasso problems. SIAM Journal on Mathematics of Data Science, 3(2):524–543, ???? 2021. CODEN SJMDAQ. ISSN 2577-0187.